OpenStudy (anonymous):

A company's revenue (in dollars) can be modeled by the function R (p) = -15p^2 + 200p +10,000 where p is the price in dollars of the company's product. What price (in dollars) will maximize the revenue? Round to the nearest penny.

OpenStudy (anonymous):

Differentiate and put equal to 0 for max

OpenStudy (anonymous):

Clarify what you mean by that...

OpenStudy (anonymous):

Differentiate R(p) with respect to p and put the resulting equal to 0 in order to calculate the value of p for which R(p) is a maximum.

OpenStudy (anonymous):

Yeah you completely lost me, aha. This is High School Math that how now been brought to another lane of confusion. I know I have to follow -b/2a... I will try to refer to another source for some help.

OpenStudy (anonymous):

It's a quadratic equation, the graph of which is a parabola. Find the vertex of the parabola.

OpenStudy (anonymous):

That in fact amounts to exactly the same thing Diff ax^2 +bx +c = 2ax + b = 0 -> x = -b/2a

OpenStudy (anonymous):

Of course it does, estudier, but if 1missmanhattan doesn't know calculus . . .

OpenStudy (anonymous):

-b/(2a) will work. Use b=200, a =-15

OpenStudy (anonymous):

Never too early to start:-)

OpenStudy (anonymous):

It's too early if the student doesn't have the requisite fundamentals. I saw a little earlier someone with a calculus question who struggled to find the area of a rectangle. It's a nice taste of things to come, though to let 1miss know that -b/2a can be found from the first derivative, but it might be better right now to relate it to something more familiar like the quadratic formula.

OpenStudy (anonymous):

Agreed, go ahead....

OpenStudy (anonymous):

Well, I think our student has already jumped ship.